# Generating optimal Gravitational-Wave template banks with metric-preserving autoencoders

### v1

@@ Abstract
We present a geometric placement algorithm for constructing template banks. We specialize in the case of Gravitational Wave searches, and use autoencoders for non-linear compression of the space of waveforms after these have been represented by a finite number of basis functions using an SVD decomposition. To ensure that the autoencoder is suitable for geometric placement we try to find a coordinate system describing the manifold of SVD coefficients such that distances in the latent and embedding space are equal. We show that the curvature of the banks is negligible and that such a system can be found. We then show that a geometric placement algorithm via a uniform grid in the latent space combined with rejection of unphysical points using a normalizing flow results in templates that, while slightly less in number than the similar construction using random forests of Ref. Wadekar:2023kym, perform slightly better in the effectualness tests, especially for high-mass binary systems. We discuss briefly how these dimensionality reduction techniques might be used in the context of cosmology, and a simple toy example where the periodicity of a flat manifold slightly complicates finding a distance-preserving coordinate system.
@@ 1 Introduction
Template banks are systematically-constructed collections of “template” waveforms or patterns used to identify or extract signals embedded in noisy data by matched filtering (comparing the data to each template). Template banks are crucial in signal-detection any field where: $(i)$ the expected signal shape is parameterizable, $(ii)$ the data is noisy, $(iii)$ direct identification is challenging. Template banks have been used in various areas of physics and astronomy, such as Gravitational Wave (GW) Astronomy (the main topic of this paper, see e.g. Wadekar:2023xbe; Abbott:2016blz; Abbott:2018gqg; LIGOScientific:2020ibl; Sachdev:2019aeb; Usman:2015kfa; Andres:2021jtr; LIGOScientific:2021djp; LIGOScientific:2024nkj; Venumadhav:2019ljl; Olsen:2022gts; Mehta:2023xbe; Chia:2023dsa; Nitz:2019gbi; Nitz:2021bwt; Nitz:2021zqy as references), Pulsar and Fast Radio Burst (FRB) Searches, Exoplanet Transit and Radial Velocity Searches, Binary Star/Compact Object Searches in X-ray and Gamma-ray Astronomy, Signal Processing in Particle Physics, Cosmic Ray and Air Shower Detection, among many more.
@@
Typically, the signal waveforms are high-dimensional but most of variation in the waveforms is along only a few effective dimensions. For example, in GW waveforms, the most variation occurs along the dimension of the binary chirp mass. The dimensionality of the template bank is set by these effective dimensions. Thus the problem becomes to capture the variation in signal waveforms into the minimum possible number of dimensions while also retaining accuracy. Apart from being of the right dimensions, the banks should have templates optimally spaced along the dimensions, otherwise this introduces redundancies (as overly-close templates give the same result) and this also increases the search computational cost. This goes under the name of geometric placement, see e.g. Roulet:2019hzy. Machine Learning architectures like autoencoders are good at dimensionality reduction but do not typically preserve the metric in the input space. We therefore construct metric-preserving autoencoders in this paper using a hybrid loss function given by $\smash{L=\alpha\times L_{\rm reconstruction}+\beta\times L_{\text{distance preservation}}}$, where $\alpha$ and $\beta$ are hyperparameters. Autoencoders are neural network architectures which have a latent bottleneck layer with lower-dimensionality than the input. The input data is therefore compressed into a latent vector and then transformed back to the original dimensions $^{\bm{c}}\rightarrow{\bm{d}}\rightarrow\hat{{\bm{c}}}$. The reconstruction loss is based on the difference between $\bm{c}$} and $\hat{{\bm{c}}}$. We will use the metric preservation loss to enforce $\smash{|{\bm{c}}-{\bm{c}}^{\prime}|\approx|{\bm{d}}-{\bm{d}}^{\prime}|}$.
@@
$\Psi_{22}({\bm{p}},f)=\langle\Psi_{22}\rangle_{\textup{sub-bank}}(f)+\sum_{A=0}^{\rm few}c^{A}({\bm{p}})\Psi^{\rm SVD}_{A}(f)\,\,$, (1)
where $\langle\cdot\rangle_{\textup{sub-bank}}$ denotes that we construct the SVD after subtracting the average over all the examples in a given sub-bank, and $\rm few$” will be equal to $9$ in the remainder of this paper. Then, the search proceeds by performing geometric placement of the templates in the much more manageable space of the SVD coefficients $c^{A}$. More precisely, the construction of the template bank involves creating a uniform grid in the space of the SVD coefficients, inspired by the fact that the mismatch between templates in a given sub-bank is well approximated by the Euclidean distance between the $c^{A}$:
@@
We then see that the distance we will want to preserve in the latent space of our autoencoder is the Euclidean distance in the space of the SVD coefficients $c^{A}$. Guidance on the dimension of the latent space comes from Ref. Wadekar:2023kym. This work performed a further dimensionality reduction after the SVD decomposition: it was observed that the $c^{A}$ lie on a $2$-dimensional sub-manifold of $\mathbb{R}^{10}$, which Ref. Wadekar:2023kym described using random forests. Hence, in this work we will autoencode the sub-manifold where the $c^{A}$ lie – $\smash{c\overset{\raisebox{-15.00002pt}{${\scriptstyle\cal E}$}}{\longrightarrow}e\overset{\raisebox{-15.00002pt}{${\scriptstyle\cal D}$}}{\longrightarrow}c}$ – and ensure both that close points in the latent space correspond to close points in the embedding space and that a uniform grid of equally-spaced $(e^{0},e^{1})$ corresponds to equally-distant $(e^{0},e^{1})$. In other words, we want to find “coordinates” $e^{\mu}$ such that the metric is flat. This is not needed if we describe the sub-manifolds using random forests as in Wadekar:2023kym, and presents an additional challenge if one wants to use autoencoders for dimensionality reduction.
The main result of the paper are Figs. 2 and 3 (that summarize the effectualness of our template banks), together with Tab. 1 (which shows the number of templates per bank and sub-bank). We compare our results using autoencoders to those of Ref. Wadekar:2023kym, showing that the template banks constructed with autoencoders perform slightly better than the ones constructed with random forests while containing a slightly smaller number of templates across all banks and sub-banks. Most importantly, however, we emphasize the difference between linear and non-linear dimensionality reduction techniques by comparing our results to what one would obtain by taking only the first $2$ SVD coefficients to construct the waveforms.
2 Template banks for Gravitational Wave searches
@@
We first review the case – already studied in Ref. Wadekar:2023kym – of using random forests to model the leading SVD coefficients $(c^{0},c^{1},\dots c^{9})$ for the phase profile of the templates in each amplitude bank and chirp mass sub-bank, comparing it to the case where we consider only the first $2$ SVD coefficients $(c^{0},c^{1})$ and put to 0 the others. Then we move to using autoencoders to model $(c^{0},c^{1},\dots c^{9})$.
@@
Our template banks are listed in Tab. 1. Our plots focus on three banks, two at the high-mass and one at the low-mass end: $\text{BBH-0},{\cal M}\in[2.6,5.3]\,{M_{\odot}}$, $\text{BBH-4},{\cal M}\in[5.8,10.8]\,{M_{\odot}}$ and $\text{BBH-12},{\cal M}\in[28.4,173.5]\,{M_{\odot}}$. In the remainder of the text and of the figures, we will denote these by “Bank 0”, “Bank 44”, “Bank 1212” for simplicity. We show the first three SVD components for Bank 0 and Bank $4$ in Fig. 1. We see that the manifold of the $c^{A}$ is very “curved”, in the sense that it is easy to see by eye that it is not a manifold described by a system of inhomogeneous linear equations. Hence we expect that nonlinear dimensionality reduction methods will perform much better than linear ones in reproducing the template waveforms.
Figure 1: The first three components of the SVD coefficients $c^{A}$ for Bank 0 (teal) and Bank 44 (orange).
2.1 Dimensionality reduction with random forests
As discussed in Section 1, in the case of the random forest our templates are constructed by first generating many examples for the phases in each bank and sub-bank, constructing the SVD decomposition as in Roulet:2019hzy, and then fitting a random forest $\rm RF$ (we use https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.RandomForestRegressor.html, with $50$ trees and maximum depth of $50$) to predict $(c^{2},c^{3},\dots c^{9})={\rm RF}(c^{0},c^{1})$ Wadekar:2023kym. In the upper-right panel of Fig. 2 we show the phase reconstruction error in the validation dataset
@@ 2.1 Dimensionality reduction with random forests
for the training of this random forest. We can compare this with the upper-left panel of that figure, where instead we take consider only the first two SVD coefficients $(c^{0},c^{1})$ and we put $(c^{2},c^{3},\dots c^{9})$ to zero.
Figure 2: Phase overlap of Eq. (3) between the validation phases and their model via random forest (upper right panel) and autoencoder (lower panel) in the selected banks $\text{BBH-0},{\cal M}\in[2.6,5.3]\,{M_{\odot}}$ (Bank 0), $\text{BBH-4},{\cal M}\in[5.8,10.8]\,{M_{\odot}}$ (Bank 44) and $\text{BBH-12},{\cal M}\in[28.4,173.5]\,{M_{\odot}}$ (Bank $12$). We see that in all three cases the random forest achieves an excellent reconstruction of the phases: less than $0.1\%$ of the validation waveforms are reconstructed with an error larger than $0.1\%$. In the autoencoder case, instead, we see that a fraction between $2\%$ and $10\%$ of validation waveforms have reconstruction error larger than $0.1\%$. The upper left panel shows instead the phase overlap between the validation phases and their reconstruction keeping only the first $2$ SVD coefficients. We see that both autoencoder and random forest greatly outperform this simple dimensionality reduction in the low-mass case. In the high-mass regime, instead, it is only the random forest that performs significantly better.
Figure 3: Effectualness of our template banks for the simple SVD dimensionality reduction $(c^{2},c^{3},\dots c^{9})=(0,0,\dots 0)$ (upper left panel), the random forest case (upper right panel) and the autoencoder case (lower panel). The full bars (dotted lines) show the result before (after) grid refinement. As discussed in the main text, for the autoencoders the grid size in the latent space is given by $\Delta e=\Delta c/{\cal N}$, where the normalization factor $\cal N$ was introduced at the stage of data preparation for the training of the autoencoders by rescaling the SVD components $c^{A}\to c^{A}/{\cal N}$. We see that generally the autoencoder performs better than the random forest in matching the test waveforms, especially after grid refinement, apart from the lowest-mass bin.
@@
Table 1: Details of the $22$ template banks. The banks are split according to $22$ amplitudes – see Fig. 4 of Wadekar:2023kym – and the sub-banks are split according to the chirp mass ($\mathcal{M}$). The lower and upper bin edges in $\mathcal{M}$ are given in the second column. From BBH-9 onwards, we only have a single sub-bank per bank. $\Delta c$ denotes the grid size in the $(c^{0},c^{1})$ space that is used to generate the templates. The last two columns show the number of templates we have according to the technique we use for the dimensionality reduction of the space of SVD coefficients. The second-to-last column shows the number of templates we have if we switch on only the 22 lowest-order SVD coefficients (“SVD22”) or use random forests to model the higher-order ones. The last column instead shows the number of templates we have if we use autoencoders with the method discussed in Section $2.$ We see that the latter results in a slightly lower number of templates than the first two.
Let us then move to the construction of the template banks themselves. [^1] The code is available at https://github.com/JayWadekar/gwIAS-HM. As in Ref. Wadekar:2023kym, we adapt the geometric placement method of Roulet:2019hzy to the case where $(c^{2},c^{3},\dots c^{9})={\rm RF}(c^{0},c^{1})$. This is straightforward: it is sufficient to construct a uniform grid in $(c^{0},c^{1})$. The grid spacing $\Delta c$ is chosen in the same way as in Roulet:2019hzy; Wadekar:2023kym, and listed in the third column of Tab. 1. In Fig. 3 we show the effectualness tests for the same three banks of Fig. 2. As in Ref. Wadekar:2023kym, for each test waveform we first unwrap its phase and take the dot product with the phase basis functions of Eq. (1) to find the $(c^{0}_{\rm test},c^{1}_{\rm test})$ coefficients of this test waveform. Then, we find the closest grid point $(c^{0}_{\rm closest},c^{1}_{\rm closest})$ in the $(c^{0},c^{1})$ space, and construct the $22$ mode using again Eq. (1) but with $(c^{0},c^{1})=(c^{0}_{\rm closest},c^{1}_{\rm closest})$ and $(c^{2},c^{3},\dots c^{9})={\rm RF}(c^{0}_{\rm closest},c^{1}_{\rm closest})$. Then, we calculate the match with the test waveform. We can then also compare this with the match obtained if we put $(c^{2},c^{3},\dots c^{9})$ to zero.
In principle, to refine this calculation, one could take the inner product of the test waveform with all the templates in the bank and find the best match. Instead of doing this, we refine our effectualness calculation by matching the test waveforms to their closest grid points on a new and refined grid with spacing $\Delta c\to\Delta c/2$. Finally, let us recall that along the $(c^{0},c^{1})$ dimensions not all the points of the rectangular grid describe physically viable waveforms. Also here, like in Wadekar:2023kym, we use the method adopted in Roulet:2019hzy of rejecting the grid points which do not have a physical waveform in their vicinity: for reference, see Fig. 4 of Roulet:2019hzy. The second column of Tab. 1 shows the number $\smash{N_{\rm templates}}$ of templates, i.e. how many of these grid points are physical for each bank and sub-bank.
In the next section we carry out the template bank construction and effectualness analysis for the same banks but using autoencoders, which will require a bit of modifications in the construction of the template grid as we hinted at in Section 1.
2.2 Dimensionality reduction with autoencoders
To start, we need to train the autoencoders to model the phases in each bank and sub-bank. We follow these steps:

### v2

@@ Abstract
Matched filtering for signal detection in noisy data requires template banks that capture variation in signal waveforms while minimizing computational cost. Dimensionality reduction of signal waveforms can be important for building efficient template banks. In various domains of physics, dimensionality reduction is very commonly performed using linear methods such as singular value decomposition (SVD). This can, however, introduce redundancies if the signals span curved, nonlinear manifolds in parameter space. Alternatively, autoencoders are a type of neural networks that can be used for non-linear dimensionality reduction. We use a variation of the autoencoder which preserves the metric in its latent space ($g_{ij}^{\text{latent}}\approx g_{ij}^{\text{physical}}$); this enables template banks to be constructed by simply placing a uniform grid in the autoencoder’s low-dimensional latent space. We apply our method for creating geometric template banks for gravitational wave searches and show that our banks require fewer dimensions compared to using the SVD basis. Our method can also be useful for other applications requiring dimensionality reduction, such as gravitational waveform modeling, fast parameter estimation and model-independent tests of general relativity. Finally, we discuss extensions to other domains including cosmological parameter estimation, and we show tests of our method in extreme cases of periodic signal manifolds.
@@ 1 Introduction
Template banks are systematically-constructed collections of “template” waveforms or patterns used to identify or extract signals embedded in noisy data by matched filtering (comparing the data to each template). Template banks are crucial in signal-detection any field where: $(i)$ the expected signal shape is parameterizable, $(ii)$ the data is noisy, $(iii)$ direct identification is challenging. Template banks have been used in various areas of physics and astronomy, such as Gravitational Wave (GW) data analysis (the main topic of this paper, see e.g. Wadekar et al. (2024a); Abbott et al. (2016, 2019, 2021); Sachdev et al. (2019); Usman et al. (2016); Andres et al. (2022); Abbott et al. (2023, 2024); Venumadhav et al. (2020); Olsen et al. (2022); Mehta et al. (2024); Chia et al. (2024); Nitz et al. (2020, 2022, 2023); Abac et al. (2025); Coogan et al. (2022) as references), Pulsar and Fast Radio Burst (FRB) searches, Exoplanet Transit and Radial Velocity searches, Binary Star/Compact Object searches in X-ray and Gamma-ray Astronomy, Signal Processing in Particle Physics, Cosmic Ray and Air Shower Detection, among many more.
Figure 1: A schematic giving an overview of our procedure for constructing the template banks using metric-preserving autoencoders.
@@
Typically, the signal waveforms are high-dimensional, but most of variation in the waveforms is effectively only along a few dimensions. For example, in GW waveforms, the dominant variation occurs along the dimension of the binary chirp mass, and the next most important variation occurs in a combination of the binary mass ratio and effective aligned spin. The dimensionality of the template bank is set by the number of such effective dimensions. Thus the problem becomes to capture the variation in signal waveforms into the minimum possible number of dimensions while also retaining accuracy. Apart from being of the right dimensions, the banks should have templates optimally spaced along the dimensions, otherwise this introduces redundancies (as overly-close templates give the same result) and this also increases the search computational cost. This approach goes under the name of geometric placement, see e.g. Roulet et al. (2019); Hanna et al. (2023); Brown et al. (2012); Phukon et al. (2025).
In various domains of physics and astronomy, dimensionality reduction is commonly performed using linear methods such as singular value decomposition (SVD) or principal component analysis (PCA). However, these methods are linear and therefore cannot capture the non-linear variation in the signal waveforms. Machine learning architectures like autoencoders are good at non-linear dimensionality reduction; autoencoders are neural network architectures which have a latent bottleneck layer with a smaller number of dimensions than the input (see Fig. 1) Hinton and Salakhutdinov (2006); Kingma and Welling (2014); Vincent et al. (2008, 2010); Bengio et al. (2007). They have been used in a few applications in astronomy and cosmology Melchior et al. (2023); Liang et al. (2023a, b); Arcelin et al. (2021). The input data is therefore compressed into a latent vector and then reconstructed back to the original dimensions $^{\bm{c}}\rightarrow{\bm{e}}\rightarrow\hat{{\bm{c}}}$. The reconstruction loss is based on the difference between $\bm{c}$} and $\hat{{\bm{c}}}$. The standard autoencoder models do not typically preserve the metric in the input space. We therefore construct metric-preserving autoencoders in this paper using a hybrid loss function given by $\smash{L=\alpha\times L_{\rm reconstruction}+\beta\times L_{\text{distance preservation}}}$, where $\alpha$ and $\beta$ are hyperparameters. We will use the metric preservation loss to enforce $\smash{|{\bm{c}}-{\bm{c}}^{\prime}|\approx|{\bm{e}}-{\bm{e}}^{\prime}|}$.
The main results of the paper are Figs. 3 and 4 (that summarize the effectualness of our template banks), together with Tab. 1 (which shows the number of templates per bank and sub-bank). We also compare our results using autoencoders to those of Ref. Wadekar et al. (2024b) who use a different ML algorithm for dimensionality reduction (random forests). We also compare the performance of the linear and non-linear dimensionality reduction techniques by showing the results of a linear model using only the first few SVD coefficients for the waveforms.
2 Geometric template banks for Gravitational Wave searches
@@
$\Psi_{22}({\bm{p}},f)=\mathrm{arg}[h({\bm{p}},f)]=\langle\Psi_{22}\rangle_{\textup{sub-bank}}(f)+\sum_{A=0}^{\rm few}c^{A}({\bm{p}})\Psi^{\rm SVD}_{A}(f)\,\,$, (1)
we only model the leading-order quadrupole mode and hence use the $22$ subscript. $\langle\cdot\rangle_{\textup{sub-bank}}$ denotes that we construct the SVD after subtracting the average over all the examples in a given sub-bank, and $\rm few$” is equal to $9$ in the remainder of this paper. Then, the search proceeds by performing geometric placement of the templates in the much more manageable space of the SVD coefficients $c^{A}$. More precisely, the construction of the template bank involves creating a uniform grid in the space of the SVD coefficients, inspired by the fact that the mismatch between templates in a given sub-bank is well approximated by the Euclidean distance between the $c^{A}$:
@@
There can however be a redundancy in the SVD coefficients, which can lead to the banks requiring more dimensions than necessary. In this work, we will therefore encode the sub-manifold where the $c^{A}$ lie ( $\smash{c\overset{\raisebox{-15.00002pt}{${\scriptstyle\cal E}$}}{\longrightarrow}e\overset{\raisebox{-15.00002pt}{${\scriptstyle\cal D}$}}{\longrightarrow}c}$ ) and ensure both that close points in the latent space correspond to close points in the embedding space and that a uniform grid of equally-spaced $(e^{0},e^{1})$ corresponds to equally-distant waveforms in the physical space. In other words, we want to find “coordinates” $e^{\mu}$ in which the overlap metric is flat.
@@
We first review the case – already studied in Ref. Wadekar et al. (2024b) – of using random forests to model the leading SVD coefficients $(c^{0},c^{1},\dots c^{9})$ for the phase profile of the templates in each amplitude bank and chirp mass sub-bank, comparing it to the case where we consider only the first $2$ SVD coefficients $(c^{0},c^{1})$ and put to 0 the others. Then we move to using autoencoders to model $(c^{0},c^{1},\dots c^{9})$.
@@
Our template banks are listed in Tab. 1. Our plots focus on three banks, two at the high-mass and one at the low-mass end: $\text{BBH-0}\,({\cal M}\in[2.6,5.3]\,{M_{\odot}}$), $\text{BBH-4}\,({\cal M}\in[5.8,10.8]\,{M_{\odot}})$ and $\text{BBH-12}\,({\cal M}\in[28.4,173.5]\,{M_{\odot}})$. In the remainder of the text and of the figures, we will denote these by “Bank 0”, “Bank 44”, “Bank 1212” for simplicity. We show the first three SVD components for Bank 0 and Bank $4$ in Fig. 2. We see that the manifold of the $c^{A}$ is not a hyperplane but rather a “curved” hypersurface (in other words, it is not a manifold described by a system of inhomogeneous linear equations). Hence we expect that nonlinear dimensionality reduction methods will perform much better than linear ones in reproducing the template waveforms.
Figure 2: We show the first three SVD coefficients $c^{A}$ obtained from decomposing the phase of gravitational wave signals: $\Psi_{22}({\bm{p}},f)=\mathrm{arg}[h({\bm{p}},f)]=\langle\Psi_{22}\rangle(f)+\sum_{A=0}^{\rm few}c^{A}({\bm{p}})\Psi^{\rm SVD}_{A}(f)$ for a range of binary parameters $\bm{p}$}. We show the coefficients for Bank 0 (teal) and Bank 44 (orange). We clearly see that the coefficients follow a narrow curved hypersurface instead of broadly filling the space. Such curved hypersurfaces cannot be compressed efficiently by linear methods like SVD. We will therefore use non-linear dimensionality reduction methods like autoencoders instead. The left and right panels are different views of the same plot.
To compare the performance of our models, we calculate the phase overlap betwen our models and the validation phases using:
@@
We show the phase overlap between the validation phases and their different models in the selected banks in Fig. 3. In the next section, we will attempt to reduce our templates to a two-dimensional model by using autoencoders. To benchmark our results, we show in the upper-left panel of that figure the phase overlap between the validation phases and their model using only the first $2$ SVD coefficients $(c^{0},c^{1})$ (while putting the rest to zero) in Eq. (1).
Figure 3: Phase overlap from Eq. (3) between the validation phases and their different models in the selected banks ${\cal M}\in[2.6,5.3]\,{M_{\odot}}$ (Bank 0), ${\cal M}\in[5.8,10.8]\,{M_{\odot}}$ (Bank 44) and ${\cal M}\in[28.4,173.5]\,{M_{\odot}}$ (Bank $12$). The upper left panel shows the phase overlap between the validation phases and their model using only the first $2$ SVD basis vectors. In the autoencoder case (lower panel), we see a much lower fraction have reconstruction error larger than $0.1\%$. We also compare our results with another machine learning algorithm, the random forest [RF] (upper right panel), where we predict the supposedly-redundant higher-order SVD coefficients $(c^{2},c^{3},\dots c^{9})$ from the first two $(c^{0},c^{1})$ Wadekar et al. (2024b). We see that both the machine learning algorithms allow for using an effectively two-dimensional parameter space and significantly outperform using just the linear SVD method, especially for low-mass banks.
Figure 4: Similar to Fig. 3, but showing the effectualness of our banks for the full GW waveform instead of just the phase overlap. We again show the performance of the two dimensional SVD model (upper left panel), autoencoder model (lower panel) and the random forest model (upper right panel). The full bars (dotted lines) show the result before (after) grid refinement (i.e., making the template grid finer). See the main text for details regarding the grid employed for constructing templates in the autoencoder latent space. We again see that both the machine learning algorithms allow for using effectively two-dimensional banks and significantly outperform using just the linear SVD method, especially for low-mass banks.
@@
Table 1: Details of the quadrupole-mode gravitational wave template banks. The banks are split according to the normalized amplitude profile – see Fig. 4 of Wadekar et al. (2024b) – and the sub-banks are split according to the chirp mass ($\mathcal{M}$). The lower and upper bin edges in $\mathcal{M}$ are given in the second column. From BBH-9 onwards, we only have a single sub-bank per bank. $\Delta c$ denotes the grid size in the $(c^{0},c^{1})$ space that is used to generate the templates. The last two columns show the number of templates we have according to the technique we use for the dimensionality reduction of the space of SVD coefficients. The second-to-last column shows the number of templates we have if we switch on only the 22 lowest-order SVD coefficients (“SVD22”) and use random forests to model the higher-order ones. The last column instead shows the number of templates we have if we use autoencoders with the method discussed in Section 3. The number of templates in the two cases is comparable, with the autoencoder resulting in a slightly lower number.
3 Dimensionality reduction with autoencoders
We train autoencoders to model the phases of waveforms in each bank and sub-bank. It is possible to use the full high-dimensional GW waveform for training the autoencoder, but it can make the training expensive. We therefore train the autoencoder on a pre-compressed version of the waveform (i.e., we use the top 10 $c^{A}$ coefficients to represent the waveform). Our goal in this paper is to probe if further compression is possible compared to just using SVD. We follow these steps for the training the autoencoder: